# revtex4-1Repair the float

### v1

@@ Abstract
We study the translocation of a flexible polymer through extended patterned pores using molecular dynamics (MD) simulations. We consider cylindrical and conical pore geometries that can be controlled by the angle of the pore apex $\alpha$. We obtained the average translocation time $\langle\tau\rangle$ for various chain lengths $N$ and the length of the pores $L_{p}$ for various values $\alpha$ and found that $\langle\tau\rangle$ scales as $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}\left(L_{p}N^{\phi}\right)$ with exponents $\gamma=3.00\pm 0.05$ and $\phi=1.50\pm 0.05$ for both patterned and unpatterned pores.
@@ I Introduction
Translocation of polymer or proteins is an ubiquitous process inside biological cells and is among one of the important processes which propel the cell functionality Alberts et al. (2002); van der Laan et al. (2007). In the last two and half decades, polymer translocation has gained considerable attention due to its technological applications such as rapid DNA sequencing, gene therapy and controlled drug delivery Salman et al. (2001); Ying et al. (2022). It is now possible to fabricate synthetic nanopores in the laboratory which can remain stable at high temperatures, external drive or change in the solvent conditions. A single strand of a biopolymer such as DNA or RNA can be electrophoretically driven through such nanopores in a controlled manner. The ion current flowing through the nanopore gets blocked when the polymer is inside the pore which gives the signature of the polymer sequence Kasianowicz et al. (1996).
@@
The translocation of polymer through nanopores depend on various parameters, such as, the length of the polymer chain, its structure, shape and the size of the pore and pore-polymer interactions Muthukumar (2007, 2011). Due to these factors, the translocation process is a complicated system dependent process. Various coarse-grained models have been used to understand the translocation phenomenon Sung and Park (1996); Muthukumar (2011); Palyulin et al. (2014); Rowghanian and Grosberg (2011); Sakaue (2007); Kantor and Kardar (2004); Dubbeldam et al. (2007); Ikonen et al. (2012a). For example, the translocation process has been studied as a one-dimensional barrier crossing problem of the translocation coordinate through a slit (i.e., a pore of unit length). In initial studies, it was assumed that the translocation time is long enough to ensure the equilibration of the polymer conformations at every stage of the translocation process Sung and Park (1996); Muthukumar (1999, 2011). Later, a majority of experimental studies with synthetic nanopores showed that the time taken by the polymer to pass the nanopore is much less than its relaxation time, indicating that the translocation is a non-equilibrium process. The tension propagation (TP) theory proposed originally by Sakaue Sakaue (2007) for an infinite chain and subsequently modified by Ikonen et al. Ikonen et al. (2012a, b); Sarabadani et al. (2014); Sarabadani and Ala-Nissila (2018) and Dubbeldam et al. Dubbeldam et al. (2012) for finite chains has proven to be successful in explaining the non-equilibrium aspect of driven polymer translocation Chen et al. (2021). The most important quantity of interest in polymer translocation studies is the dependence of the average translocation time, $\langle\tau\rangle$, on the chain length $N$. Numerical simulations and experiments have indicated that $\langle\tau\rangle\sim N^{\beta}$. Different values of $\beta$ have been observed in different studies, suggesting that $\langle\tau\rangle$ depends on various physical parameters Milchev (2011). For a self-avoiding flexible chain, $\langle\tau\rangle$ is known to scale as $\langle\tau\rangle\sim N^{1+\nu}$, here $\nu=3/4$ in two dimensions Chuang et al. (2001); Luo et al. (2007); Sarabadani and Ala-Nissila (2018), and $\nu\approx 0.59$ in three dimensions Doi et al. (1988); Chuang et al. (2001); Kantor and Kardar (2004).
@@
Experimental studies on polymer translocation involve pores of finite lengths in which pore-polymer interactions play a dominant role Storm et al. (2003); Wanunu and Meller (2007); Chen et al. (2004); Lan et al. (2011); Liu et al. (2012). There have been theoretical and computational studies on finite-length pores where the shape of the extended pore is cylindrical with different patterned pore-polymer interactions Luo et al. (2007, 2008); Cohen et al. (2012a, b); Katkar and Muthukumar (2018); Kumar et al. (2018), or conical Nikoofard et al. (2013, 2014); Nikoofard and Fazli (2015); Sharma et al. (2022); Sharma (2024); Upadhyay et al. (2024a, b). For cylindrical pores, it was estimated that for a fixed external drive $\langle\tau\rangle$ depends on the patterning of the pores. It was found that $\langle\tau\rangle$ is the maximum for a fully attractive pore and is minimum for a pore having repulsive exit Cohen et al. (2012b); Kumar et al. (2018). It was argued that the pattern stickiness of extended pores can be exploited to predict heteropolymer sequences with different bending rigidities to a high degree of accuracy by translocating them through multiple pores Kumar et al. (2018). For conical pores, we have one more parameter–the cone apex angle, which we represent by $\alpha$. For $\alpha=0$, we have a cylindrical pore. For a conical pore with attractive pore-polymer interactions, $\langle\tau\rangle$ was found to be a nonmonotonic function of $\alpha$ for lower driving forces, which becomes monotonic for higher forces Sharma et al. (2022); Sharma (2024). The length of the extended pore, $L_{p}$, introduces an additional length scale to the problem. Although there have been many scaling studies of $\langle\tau\rangle$ on the length of the polymer $N$, not much is discussed in the literature on the dependence of $\langle\tau\rangle$ on $L_{p}$. We try to address this in this paper.
@@
In this paper, we extend our work on conical pores Sharma et al. (2022); Sharma (2024) to study the effect of different pore patterns on the driven-translocation process. The patterning is created along the pore length by having various sections with attractive and repulsive interactions. The entrance of the pores is always attractive, so that the polymer can enter the pore with ease. The exit from the pore can either be attractive or repulsive. The presence of an external drive along the channel often facilitates the translocation process. We present our study on translocation behavior by varying the pore length $L_{p}$ and the polymer length $N$. The length of the pore is increased in such a manner that the pore-patterning ratios remain conserved. We record the translocation time through such pores for different external drives.
@@
The paper is organized as follows. In Sec. II, we define our model and simulation details. In Sec. III, we study the effect of the pore length, $L_{p}$, on $\langle\tau\rangle$. We also discuss the effect of pattered pores on $\langle\tau\rangle$ and establish scaling relations between $\langle\tau\rangle$, $L_{p}$ and the polymer length $N$ for extended pores. In Sec. IV, we conclude by providing a discussion of the obtained scaling exponents.
@@
Figure 1: Top row shows a schematic diagram of pore-polymer set up for translocation in two-dimensions. The pore is 5 beads long with half apex angle $\alpha$. The pore diameter on the cis end is $d=2.25\sigma$, whereas it is $y=d+2L_{p}\tan\alpha$ on the trans end. The flexible polymer is made up of beads of diameter $\sigma=1$ connected with springs. The subsequent rows show patterned pores with different pore lengths, $L_{p}=$ 5, 10, and 15, respectively. Blue beads are attractive while yellow and green beads are repulsive in interaction with polymer chain.
@@ II Model
We study the pore-polymer system in two-dimensions. The polymer is modeled as a bead-spring chain Grest and Kremer (1986), where the consecutive beads of diameter $\sigma$ are connected with a spring of spring constant $K$ interacting via a harmonic potential given by:
@@
where, $r_{0}$ is the equilibrium distance. The non-bonded polymer beads interact via Weeks-Chandler-Andersen (WCA) potential $U_{\textrm{nb}}(r)$ of the form:
@@
where, $\epsilon$ is the strength of the potential. The cutoff distance $r_{\rm min}=2^{1/6}\sigma$ is set at the minimum of the potential. The pores are static and are constructed with coarse-grained beads of size $\sigma$. We study three different types of extended pores as considered in Ref. Kumar et al. (2018):
@@
    Pore A is an attractive pore. Polymer beads interact with pore beads via the attractive Lennard Jones (LJ) potential, $U_{\textrm{pp}}(r)$:
@@
    where, $\epsilon_{p}$ is the attractive strength of the pore and $r_{c}=2.5\sigma$ is the cut-off distance.
@@
    Pore C has an attractive entrance and a repulsive exit. The polymer beads interact with the beads of the attractive section of the pore via the LJ potential and the beads of the repulsive section via WCA potential.
@@
In addition to the above potentials, an external pulling force, $f_{ext}$, which mimics an electrophoretic force, is applied inside the pore in the positive $x$-direction to ensure that the polymer translocates through the pore from the cis to the trans side. The force $f_{ext}$ varies along the pore extension as
$f_{ext}=\frac{f_{0}d}{(d+2x\tan\alpha)},$ (4)
@@
here $f_{0}$ is the magnitude of the applied force at the pore opening $d$, which is equal to the constant force inside a linear channel with $\alpha=0^{\circ}$. Therefore, for a conical pore, the force decreases along its length. The schematic diagram of our coarse-grained model is shown in Fig. 1 (top row). The subsequent rows show the schematic diagram of three different pore patterns that we will study in this paper. The second row shows three different patterned pores for length $L_{p}=5$, here the pore in Fig. 1(a1) is a completely attractive pore (Pore A), while the pore in Fig. 1(b1) has an attractive entrance and an attractive exit but embedded with a repulsive bead in the middle (Pore B), and the pore in Fig. 1(c1) has an attractive entrance but a repulsive exit (Pore C). In the third and fourth rows, we show pores for lengths $L_{p}=10$, and $L_{p}=15$, respectively. The length of the pore is increased in such a manner that the ratio of pore patterning remains conserved.
@@
The dynamics of the polymer is governed by the stochastic Langevin equation. The equation of motion for the position of $i$th monomer is given by
$m\ddot{\vec{r_{i}}}=-\eta\dot{\vec{r_{i}}}-\nabla\sum{(U_{bond}+U_{nb}+U_{pp})}+f_{ext}+\zeta,$ (5)
where $m$ and $r_{i}$ are the mass and the position of the $i^{th}$ bead of the polymer, $\eta$ is the friction coefficient of the fluid, $\zeta$ is an uncorrelated random noise that obeys the fluctuation-dissipation relation $\langle\zeta_{i}(t)\zeta_{j}(t^{\prime})\rangle=2K_{B}T\eta\delta_{ij}\delta(t-t^{\prime})$. In our model, $\epsilon$, $\sigma$ and $m$ set the units of energy, length, and mass, respectively. This sets the unit of time as $(m\sigma^{2}/\epsilon)^{1/2}$. The force value is set in the unit of $\epsilon/\sigma$ and the spring constant in the units of $\epsilon/\sigma^{2}$. Using this unit, the dimensionless parameters $k_{B}T=1$, $\eta=1$, $d=2.25$, and $K=500$ are chosen in the simulations. For every simulation data point, the results are averaged over 1500-2000 successful translocation events. MD trajectories are obtained using LAMMPSThompson et al. (2022) software.
@@ III.1 Effect of pore length
Figure 4: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for an extended pore ($\alpha=0$) for three different driving forces (a) 0.2, (b) 0.6, and (c) 1.0. The lines are the scaling function $\mathcal{F}(x)\sim 1/x^{p}$ with exponent $p=0.85$.
@@
The length of the pores, $L_{p}$, plays an important role in the translocation of a polymer through a membrane channel. To study how the translocation time depends on the length of the pores, we consider a polymer of length $N=128$ and translocate it through pores of varying lengths $L_{p}\in[5\sigma$, $10\sigma]$. For each pore length, we also vary the angle of the cone apex, $\alpha$, in the range $\alpha\in[0,5]$ degrees. The magnitude of the external force is kept constant at $f_{0}=1$. The pore-polymer interaction is attractive in nature and is modeled using Eq.(3). In Fig. 2, we have plotted the average translocation time $\langle\tau\rangle$ as a function of $\alpha$ for various $L_{p}$. The plot clearly shows that, irrespective of the value of $\alpha$, $\langle\tau\rangle$ decreases with increasing pore length $L_{p}$.
For $\alpha=0^{\circ}$ the value of $\langle\tau\rangle$ obtained for $L_{p}=5\sigma$ is almost three times that for $L_{p}=10\sigma$. This can be understood as follows: At $\alpha=0^{\circ}$ the pore is cylindrical in geometry with a diameter $d=2.25\sigma$, which allows only a single-file entry and exit from the pore. For such symmetrical pores, the extent of the external force increases as the pore length is increased. For example, for a completely filled pore of length $L_{p}=10\sigma$, the number of beads experiencing force is almost twice the number of beads for that of $L_{p}=5\sigma$. Therefore, the polymer can be driven easily through a longer pore.
For $\alpha>0^{\circ}$, i.e., for conical pores, the diameter of the channel increases as one moves from the apex to the base of the pore. The diameter at the narrow entrance remains the same as that for a cylindrical pore (i.e., $d=2.25\sigma$), while the diameter of the wider end increases by an amount $2L_{p}\sin\alpha$ [i.e., $d(x)=d+2x\sin\alpha$]. Now as the area available on the wider end of the conical channel increases, the single-file nature of the translocation process is no longer true and more configurations of the polymer are possible, this gain in the conformational entropic force Nikoofard et al. (2013) assists the polymer to escape faster from the channel. This is indeed seen for shorter pore lengths, e.g., $L_{p}=5$ and 6, where $\langle\tau\rangle$ decreases on increasing $\alpha$ and saturates for $\alpha>2^{\circ}$. In contrast, for longer pore lengths, we see an opposite behavior. As the length of the pore increases, the diameter at the wider end increases proportionately. Since the external force inside the channel decreases as $f(x)\sim 1/x$ (see Eq. (4)), it also becomes weaker at the wider pore exit for longer pores. Further, since the pore is attractive in nature there is a higher possibility for the polymer to have a configuration where the two consecutive monomers of the chain occupy side-by-side position along the $y$-direction at some location $x$ in the channel, thus hindering translocation. Combining all these factors, it was observed that the total translocation time for longer pores increases as the pore apex angle $\alpha$ is increased (e.g., $L_{p}=10\sigma$ in Fig. 2).
Figure 5: Comparison of $\tau$ vs. $\alpha$ for three different pore sizes $L_{p}=$ 5, 10 and 15 for two different forces ($f_{0}=$) 0.6 and 1.0. The length(NN) of the flexible polymer chain is 128 beads.

### v2

# Polymer translocation through extended patterned pores in two dimensions: Scaling of the total translocation time

@@ Abstract
We study the translocation of a flexible polymer through extended patterned pores using molecular dynamics (MD) simulations. We consider cylindrical and conical pore geometries that can be controlled by the angle of the pore apex $\alpha$. We obtained the average translocation time $\langle\tau\rangle$ for various chain lengths $N$ and the length of the pores $L_{p}$ for various values $\alpha$ and found that $\langle\tau\rangle$ scales as $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}\left(L_{p}N^{\phi}\right)$ with exponents $\gamma=3.00\pm 0.05$ and $\phi=1.50\pm 0.05$ for both patterned and unpatterned pores, respectively.
@@ I Introduction
Polymer and protein translocation through narrow pores is a fundamental process in biology and an emerging paradigm for nanoscale transport in technology. Translocation of polymer or proteins is an ubiquitous process inside biological cells and is among one of the important processes that propel the cell functionality Alberts et al. (2002); van der Laan et al. (2007). In the past two and half decades, polymer translocation has gained considerable attention due to its technological applications such as rapid DNA sequencing, gene therapy and controlled drug delivery Salman et al. (2001); Ying et al. (2022). Recent advances in nanofabrication have enabled stable synthetic nanopores that can operate under high voltages, elevated temperatures, and various solvent conditions. These pores have become controllable experimental systems to explore the physics of single‐molecule transport. A single strand of a biopolymer such as DNA or RNA can be electrophoretically driven through such nanopores in a controlled manner. The ion current flowing through the nanopore gets blocked when the polymer is inside the pore which gives the signature of the polymer sequence Kasianowicz et al. (1996).
@@
The translocation of polymer through nanopores depends on various parameters, such as, the length of the polymer chain, its structure, shape and the size of the pore and pore-polymer interactions Muthukumar (2007, 2011). Consequently, polymer translocation is a complex, system-dependent nonequilibrium process whose understanding requires both theoretical modelling and computational simulations. Various coarse-grained models have been used to understand the translocation phenomenon Sung and Park (1996); Muthukumar (2011); Palyulin et al. (2014); Rowghanian and Grosberg (2011); Sakaue (2007); Kantor and Kardar (2004); Dubbeldam et al. (2007); Ikonen et al. (2012a). For example, the translocation process has been studied as a one-dimensional barrier crossing problem of the translocation coordinate through a slit (i.e., a pore of unit length). In initial studies, it was assumed that the translocation time is long enough to ensure the equilibration of the polymer conformations at every stage of the translocation process Sung and Park (1996); Muthukumar (1999, 2011). Later, a majority of experimental studies with synthetic nanopores showed that the time taken by the polymer to pass the nanopore is much less than its relaxation time, indicating that translocation is a non-equilibrium process. The tension propagation (TP) theory proposed originally by Sakaue Sakaue (2007) for an infinite chain and subsequently modified by Ikonen et al. Ikonen et al. (2012a, b); Sarabadani et al. (2014); Sarabadani and Ala-Nissila (2018) and Dubbeldam et al. Dubbeldam et al. (2012) for finite chains has proven to be successful in explaining the non-equilibrium aspect of driven polymer translocation Chen et al. (2021). The most important quantity of interest in polymer translocation studies is the dependence of the average translocation time, $\langle\tau\rangle$, on the chain length $N$. Numerical simulations and experiments have indicated that $\langle\tau\rangle\sim N^{\beta}$. Different values of $\beta$ have been observed in different studies, suggesting that $\langle\tau\rangle$ depends on various physical parameters Milchev (2011). For a self-avoiding flexible chain, $\langle\tau\rangle$ is known to scale as $\langle\tau\rangle\sim N^{1+\nu}$, here $\nu=3/4$ in two dimensions Chuang et al. (2001); Luo et al. (2007); Sarabadani and Ala-Nissila (2018), and $\nu\approx 0.59$ in three dimensions Doi et al. (1988); Chuang et al. (2001); Kantor and Kardar (2004).
@@
While early theoretical work considered a very short or unit‐length pore, most realistic nanopores are extended and of finite length, often with diameters of tens of monomer long. In such pores, the details of the pore–polymer interaction and spatial patterning along the pore axis strongly influence the dynamics Storm et al. (2003); Wanunu and Meller (2007); Chen et al. (2004); Lan et al. (2011); Liu et al. (2012). There have been theoretical and computational studies on finite-length pores where the shape of the extended pore is cylindrical, with different patterned pore-polymer interactions Luo et al. (2007, 2008); Cohen et al. (2012a, b); Katkar and Muthukumar (2018); Kumar et al. (2018), or conical Nikoofard et al. (2013, 2014); Nikoofard and Fazli (2015); Sharma et al. (2022); Sharma (2024); Upadhyay et al. (2024a, b). For cylindrical pores, it was estimated that for a fixed external drive $\langle\tau\rangle$ depends on the patterning of the pores. It was found that $\langle\tau\rangle$ is the maximum for a fully attractive pore and is the minimum for a pore having a repulsive exit Cohen et al. (2012b); Kumar et al. (2018). It was argued that the pattern stickiness of extended pores can be exploited to predict heteropolymer sequences with different bending rigidities to a high degree of accuracy by translocating them through multiple pores Kumar et al. (2018). For conical pores, we have one more parameter–the cone apex angle, which we represent by $\alpha$. For $\alpha=0$, we have a cylindrical pore. For a conical pore with attractive pore-polymer interactions, $\langle\tau\rangle$ was found to be a nonmonotonic function of $\alpha$ for lower driving forces, which becomes monotonic for higher forces Sharma et al. (2022); Sharma (2024).
The length of the extended pore, $L_{p}$, introduces a second independent length scale in addition to the polymer contour length $N$. Understanding how these two compete under different interaction patterns is essential for controlling transport through synthetic or biological nanopores. Although there have been many scaling studies of $\langle\tau\rangle$ on the length of the polymer $N$, not much is discussed in the literature on the dependence of $\langle\tau\rangle$ on $L_{p}$. We try to address this in this paper.
@@
In this work we use two-dimensional Langevin dynamics simulations to study the driven translocation of a flexible polymer through extended conical pores having different attractive–repulsive patterns. The two-dimensional approach enables long-chain statistics and extensive averaging, while retaining the essential physics of polymer stretching and confinement. We extend our work on conical pores Sharma et al. (2022); Sharma (2024) to study the effect of different pore patterns on the driven-translocation process. The patterning is created along the pore length by having various sections with attractive and repulsive interactions. The entrance of the pores is always attractive, so that the polymer can enter the pore with ease. The exit from the pore can either be attractive or repulsive. The presence of an external drive along the channel often facilitates the translocation process. We present our study on translocation behavior by varying the pore length $L_{p}$ and the polymer length $N$. The key questions we address are: (i) how does the mean translocation time depend simultaneously on $N$ and $L_{p}$, and (ii) whether distinct pore patterns produce universal or pattern-specific scaling behavior.
@@
The rest of this paper is organized as follows. In Sec. II, we define our model and simulation details. In Sec. III, we study the effect of the pore length, $L_{p}$, and chain length, $N$, on $\langle\tau\rangle$. We also discuss the effect of pattered pores on $\langle\tau\rangle$ and establish scaling relations between $\langle\tau\rangle$, $L_{p}$ and the polymer length $N$ for extended pores. In Sec. IV, we conclude by providing a discussion of the obtained scaling exponents.
@@
Figure 1: Top row shows a schematic diagram of pore-polymer set up for translocation in two-dimensions. The pore is 5 beads long with a half apex angle $\alpha$. The pore diameter on the cis end is $d=2.25\sigma$, whereas it is $y=d+2L_{p}\tan\alpha$ on the trans end. The flexible polymer is made up of beads with diameter $\sigma=1$ connected with springs. The subsequent rows show patterned pores with different pore lengths, $L_{p}=$ 5, 10, and 15, respectively. The blue beads are attractive while the yellow and green beads are repulsive in interaction with polymer chain.
@@ II Model
All simulations are performed in two dimensions (2D). The reduced dimensionality allows extensive statistical sampling and direct comparison with earlier 2D translocation studies Chuang et al. (2001); Luo et al. (2007). The essential physics of polymer stretching and confinement within narrow pores remains captured in 2D while significantly lowering computational cost.
The polymer is modeled as a bead-spring chain Grest and Kremer (1986), where the consecutive beads with diameter $\sigma$ are connected with a spring of spring constant $K$ interacting via a harmonic potential given by:
@@
where $r_{0}=1.0\sigma$ is the equilibrium bond length. The non-bonded polymer beads interact via the Weeks-Chandler-Andersen (WCA) potential $U_{\textrm{nb}}(r)$ of the form:
@@
where $\epsilon$ is the strength of the potential. The cutoff distance $r_{\rm min}=2^{1/6}\sigma$ is set at the minimum of the potential. Unless otherwise stated, we set $\epsilon=1.0$ and use the same energy scale for all interactions.
The pores are static and are constructed with coarse-grained beads of size $\sigma$. We study three different types of extended pores as considered in Ref. Kumar et al. (2018):
@@
    Pore A is an attractive pore. Polymer beads interact with pore beads via the attractive Lennard-Jones (LJ) potential, $U_{\textrm{pp}}(r)$:
@@
    where $\epsilon_{p}=1.5\epsilon$ is the attractive strength of the pore and $r_{c}=2.5\sigma$ is the cut-off distance.
@@
    Pore C has an attractive entrance and a repulsive exit. The polymer beads interact with the beads of the attractive section of the pore via the LJ potential and the beads of the repulsive section via the WCA potential.
The entrance of all pores is kept attractive to ensure capture of the polymer, while the exit region determines the ease of release. The total pore pattern is scaled proportionally when the pore length $L_{p}$ is varied so that the relative fraction of attractive and repulsive segments remains constant.
@@
In addition to the above potentials, an external pulling force, $\vec{f}_{\textrm{ext}}$, which mimics an electrophoretic drive, is applied inside the pore in the positive $x$-direction to ensure that the polymer translocates through the pore from the cis to the trans side. The force magnitude varies along the pore extension as
$f_{\textrm{ext}}(x)=\frac{f_{0}d}{(d+2x\tan\alpha)},$ (4)
@@
where $f_{0}$ is the driving force at the pore entrance of width $d$. For a conical pore ($\alpha>0$), this form maintains a constant total integrated force along the pore length, causing the local field to decrease gradually from the entrance to the exit. For $\alpha=0^{\circ}$, the same equation reduces to a constant field in a cylindrical pore, as expected. The schematic diagram of our coarse-grained model is shown in Fig. 1 (top row). The subsequent rows show the schematic diagram of three different pore patterns that we study in this paper. The second row shows three different patterned pores for length $L_{p}=5$: Fig. 1(a1) corresponds to a fully attractive pore (Pore A), Fig. 1(b1) shows a pore with attractive entrances and exits but a repulsive middle segment (Pore B), and Fig. 1(c1) shows a pore with an attractive entrance and a repulsive exit (Pore C). In the third and fourth rows, we show pores for lengths $L_{p}=10$ and $L_{p}=15$, respectively.
@@
The dynamics of the polymer is governed by the stochastic Langevin equation. The equation of motion for the position of the $i$th monomer is
$m\ddot{\vec{r_{i}}}=-\eta\dot{\vec{r_{i}}}-\nabla_{i}(U_{\textrm{bond}}+U_{\textrm{nb}}+U_{\textrm{pp}})+\vec{f}_{\textrm{ext}}+\vec{\zeta_{i}},$ (5)
where $m$ and $\vec{r_{i}}$ are the mass and position of the $i$th bead, respectively, $\eta$ is the friction coefficient of the fluid, and $\vec{\zeta_{i}}$ is a random Gaussian noise obeying the fluctuation–dissipation relation $\langle\zeta_{i\mu}(t)\zeta_{j\nu}(t^{\prime})\rangle=2k_{B}T\eta\delta_{ij}\delta_{\mu\nu}\delta(t-t^{\prime})$.
In our model, $\epsilon$, $\sigma$ and $m$ set the units of energy, length, and mass, respectively. This sets the unit of time as $(m\sigma^{2}/\epsilon)^{1/2}$. The force value is set in units of $\epsilon/\sigma$ and the spring constant in units of $\epsilon/\sigma^{2}$. Using this system of units, the dimensionless parameters $k_{B}T=1$, $\eta=1$, $d=2.25$, and $K=500$ are chosen in the simulations. The choices of $\eta$ and $d$ follow standard coarse-grained studies Luo et al. (2008); Cohen et al. (2012b), ensuring realistic frictional damping and pore geometry comparable with experimental nanopores.
For every simulation data point, the results are averaged over 1500–2000 successful translocation events. Error bars representing one standard error of the mean are shown in all plots. Molecular-dynamics trajectories are obtained using LAMMPS Thompson et al. (2022) software.
@@ III.1 Effect of pore length
Figure 4: $\langle\tau\rangle/N^{\gamma}$ as a function of $L_{p}N^{\phi}$ for an extended pore ($\alpha=0$) for three different driving forces (a) 0.2, (b) 0.6, and (c) 1.0. The lines are the scaling function $\mathcal{F}(x)\sim 1/x^{p}$ with exponent $p=0.85$.The dashed line above (below) is for exponent value p=0.80(p=0.90)p=0.80\ (p=0.90)
@@
The length of the pores, $L_{p}$, plays an important role in the translocation of a polymer through a membrane channel. To study how the translocation time depends on the length of the pores, we consider a polymer of length $N=128$ and translocate it through pores of varying lengths $L_{p}\in[5\sigma,10\sigma]$. (Error bars in all figures denote one standard error based on 1500–2000 successful events.) For each pore length, we also vary the angle of the cone apex, $\alpha$, in the range $\alpha\in[0,5]$ degrees. The magnitude of the external force is kept constant at $f_{0}=1$. The pore-polymer interaction is attractive in nature and is modeled using Eq.(3). In Fig. 2, we have plotted the average translocation time $\langle\tau\rangle$ as a function of $\alpha$ for various $L_{p}$.
Figure 2 shows that (i) for cylindrical pores ($\alpha=0^{\circ}$) the mean translocation time $\langle\tau\rangle$ decreases strongly with pore length $L_{p}$, and (ii) for conical pores ($\alpha>0$) $\langle\tau\rangle$ decreases with $\alpha$ for short pores but increases with $\alpha$ for long pores. This behavior can be understood as a competition between two effects: (a) the widening exit increases the available configuration space for the leading segment and reduces confinement-related hindrance near the exit (entropic relief), which facilitates the final emptying for short pores; and (b) the imposed driving field in a cone decreases along the pore, $f_{\rm ext}(x)=f_{0}d/(d+2x\tan\alpha)$ (Eq. (4)), so an extended weak-drive region develops near the exit, making long conical pores bottleneck-limited. Because polymer translocation is a nonequilibrium process involving tension propagation and conformation changes along the chain, we do not attempt a point-particle drift-time theory; instead, the finite-size scaling analysis presented below provides the quantitative characterization of $\langle\tau\rangle$ across $(N,L_{p},\alpha)$ and pore patterning.
Role of attractive interactions. In short cones, the entropic assistance dominates and $\langle\tau\rangle$ decreases with $\alpha$ before saturating once the exit is sufficiently wide. In long cones, the weakening of the local drive near the exit, outweighs the entropic gain; therefore, $\langle\tau\rangle$ increases with $\alpha$. Attractive pore–polymer interactions further amplify this effect by increasing residence near the walls and promoting side-by-side monomer arrangements that impede axial progress, again more relevant at larger $L_{p}$ where the low-drive region is longer.
Consistency with the finite-size scaling analysis. The same geometry–drive competition is reflected in the scaling form $\langle\tau\rangle\sim N^{\gamma}\mathcal{F}(L_{p}N^{\phi})$ with $\mathcal{F}(x)\sim x^{-p}$: for $\alpha=0$, increasing $L_{p}$ increases the number of driven monomers and decreases $\langle\tau\rangle$ (smaller $\mathcal{F}$); for $\alpha>0$, the weakening of the local drive near the exit with increasing $L_{p}\tan\alpha$ counteracts and eventually reverses the entropic speed-up, consistent with the crossover observed in Fig. 2. The measured exponents also satisfy $\beta\approx\gamma-(\phi p)/2$, linking the chain-length and pore-length dependence.
III.2 Effect of chain length NN